OpenStudy (danise):

Can someone teach me properties of exponents? I don't understand what to do or how the step begin. the 7 rules don't make sense to me how are they suppose to help me and are you only suppose to simplify??? I'm so confused The only one I understand is this: (h^7k^3)0 and you used the rule x^0=1 so the answer is 1

OpenStudy (anonymous):

Do you have any specific problems you need help with?

OpenStudy (anonymous):

If you are multiplying the same bases with different exponents, you add them. If you are dividing the same bases with different exponents, you subtract them. If your exponent is negative, that is saying the variable(or base) with that exponent is in the wrong place and needs to be on the opposite side of the fraction. (if the exponent is negative in numerator, move it to the denominator to make it positive; if the exponent is negative in denominator, move it to numerator to make it positive)

OpenStudy (danise):

It not letting me post the photo

OpenStudy (anonymous):

write one of them using the equation button for posting a comment

OpenStudy (danise):

(2x^3y^7)^-2

OpenStudy (anonymous):

okay you aare raising exponents to another exponent, you have to distribute that -2 outside exponent to multiply by each exponent within the parentheses. don't forget that the 2 on the inside parentheses has an invisible/understood exponent of 1

OpenStudy (anonymous):

\[(2^{1\times -2}x ^{3\times-2}y^{7\times-2})\]

OpenStudy (anonymous):

Those 'x's are multiplication symbols

OpenStudy (danise):

Ok

OpenStudy (anonymous):

So then you will need to simplify, and get rid of any negative exponents by moving them to the denominator

OpenStudy (danise):

That's it ???

OpenStudy (anonymous):

Yep! If you just remember the rules, you'll be golden :)